OpenStudy (anonymous):

How can I simplify the radical 4th root of g^3 h^5 over 9r^6

OpenStudy (anonymous):

It could be rewritten as this: \[(g^{1/4}h^{5/4})/(9r^6)\] I guess that would be simplification. The original looked like this, right? \[(\sqrt[4]{g^3h^5})/(9r^6)\]

OpenStudy (anonymous):

@chol ?

OpenStudy (anonymous):

Yeah, yeah

OpenStudy (anonymous):

was this helpful? What I did was rewrite it as (g^3 h^5) to the power of 1/4 instead of using fourth degree radical.

OpenStudy (anonymous):

\[\sqrt[4]{g^3h^5 \over 9r^6}\] It's like this

OpenStudy (anonymous):

OOHHH... ok, then its a little different. I didn't realize the radical was over all of it. Give me 2 minutes:)

OpenStudy (anonymous):

All right, this is what I have...

OpenStudy (anonymous):

\[(g^{3/4}h^{5/4})/(\sqrt{3}r^{3/4})\] Basically, I did the same thing as I did in the other one.

OpenStudy (anonymous):

@chol ? you there? lol...

OpenStudy (anonymous):

Yes I am

OpenStudy (anonymous):

I find it hard in the radical. What will I do with it?

OpenStudy (anonymous):

You can't do anything with the radical 3. You'll have to keep it, but it's simplified because the radical is no longer over the whole thing and there are no fourth roots.

OpenStudy (anonymous):

So that's it? How about the exponents?

OpenStudy (anonymous):

You have to have exponents. They are the exponents of variables, not numbers, so you can't get rid of them. What I gave you, believe or not, is as simplified as it gets:D

OpenStudy (anonymous):

So rationalizing the denominator is not needed?

OpenStudy (anonymous):

Not the way I was taught... o.O In this case, it could make the equation more complicated.

OpenStudy (anonymous):

You're right. But in the examples of our professor, he rationalizes the denominator even if it gets even more complicated as long as you remove the radicals. And it's really mind-shaking.

OpenStudy (anonymous):

my god... you think you can rationalize it yourself? I'm a little short on time:/ Hope I was helpful! Good luck with your studies:)

OpenStudy (anonymous):

I don't know. But I'll try. Thank you a lot.

OpenStudy (anonymous):

no problem:)